1a. Suppose a hydrogen-oxygen fuel-cell generator was used to produce electricity for a house. Use the balanced redox reactions and the standard cell potential to predict the volume of hydrogen gas (at STP) required each month to generate the electricity needed for a typical house. Assume the home uses 1400 kWh of electricity per month.

1b. Suppose a hydrogen-oxygen fuel-cell generator was used to produce electricity for a house. Use the balanced redox reactions and the standard cell potential to predict the volume of hydrogen gas (at STP) required each month to generate the electricity needed for a typical house. Assume the home uses 1600 kWh of electricity per month.
Answer #1

Solution :- Given that : cas tuh electvitity per month = 1400 -The half cell reactiong in a fuel.cell: Hz. -> 2H+ + Qe o Fo00» 5.04 X10 9 I al Energy 1 mol of H2 = above lat stp. Volume of H, 9ae - 5.4 Mo1 ха хәә-4 HTH 780 Qa5,792,000,000 471780 - 4.-Total = 31x2u = 744 hours on a month. 1600 knett = 1600x3.6 *106 = 5.76X1095 AG -DFE Oh x 96.500 X123 = 474780 3/a modes of